# Infrared-Assisted Single-Stage Framework for Joint Restoration and Fusion of Visible and Infrared Images under Hazy Conditions

### v1

@@ 4.2 Comparison with State-of-the-art Methods
There are two existing methodologies for fusing hazy visible images with infrared images. The first methodology involves initially applying advanced image dehazing algorithms to remove haze from the visible images, followed by fusing the dehazed images with the infrared images. For this purpose, we select the latest and most effective dehazing methods, specifically DIACMP [45] and Dehazeformer [30]. Next, we apply representative IR-VIS image fusion methods, such as MLFusion [13], U2Fusion [36], LRRNet [15], ALFusion [17], TIMFusion [24], and MRFS [44], to fuse the dehazed visible images with the infrared images. The second methodology employs the Text-IF method [40], which directly restores and fuses hazy images with the assistance of text information.
@@
Table 1: Quantitative results of six ablation experiments on the MSRS dataset, with the values in red indicating the best performance and those in blue indicating the second-best.
@@
Figure 3 sequentially displays the visual effects of fusion results obtained using different methods on the MSRS, M3FD, and RoadScene datasets. As shown in Figure 3, our method not only effectively removes haze but also successfully integrates complementary information from the IR-VIS images, producing clear and coherent fusion results. Furthermore, our approach efficiently restores brightness information, enhancing overall contrast while avoiding the introduction of noise into the image. In contrast, alternative methods, which treat dehazing and fusion as independent processes, do not facilitate effective fusion during the dehazing stage, resulting in lower-quality fusion outcomes. When faced with complex weather conditions, traditional dehazing methods often struggle to accurately restore sky regions. However, our proposed end-to-end joint framework, leveraging infrared information, effectively helps the network restore these sky regions, thereby achieving superior restoration and fusion results. Additionally, while the Text-IF can perform both dehazing and fusion, it does not exhibit significant advantages in either task.
To further substantiate the advantages of our method, we present violin plots that quantitatively compare our approach with traditional methods. The detailed data comparison can be found in the supplementary materials. As illustrated in the figure 4, our method consistently outperforms others across three different datasets and five evaluation metrics, demonstrating a more concentrated data distribution, which indicates more robust and stable performance in hazy image fusion. Moreover, our method achieves the best average performance among all methods in terms of the $Q_{PI}$, $Q_{NIQE}$, and $Q_{SF}$, further affirming the exceptional restoration capabilities of our model.
4.3 Ablation Study
We design six experimental settings to evaluate the effectiveness of each module. In the first setting, we remove the PGM-generated prompt ${\bm{\hat{P}}}_{ir}$, the HDE module, and the process of supplementing visible features with infrared features based on haze density, directly inputting ${\bm{F}}_{vi}$ into the Transformer Block for dehazing (denoted as “w/o ${\bm{F}}_{ir}$”) to assess the auxiliary role of infrared information. In the second setting, we replace the operation in Eq. (5) by directly adding ${\bm{\hat{F}}}_{ir}$ and ${\bm{F}}_{vi}$ (denoted as “w/o HDE”) to validate the impact of haze-density-based infrared integration. In the third setting, we remove the prompt embedding for ${\bm{\hat{P}}}_{ir}$ and the PEB module, injecting infrared features extracted by the encoder, guided by haze density ${\bm{H}}$, directly into ${\bm{F}}_{vi}$ (denoted as “w/o ${\bm{\hat{P}}}_{ir}$”) to evaluate the effect of prompt embedding. In the fourth setting, we test the effectiveness of the PEB module by removing it from IA-FRM, denoted as “w/o FR-PEB”. In the fifth setting, we remove ${\bm{\hat{P}}}_{vi}$ from MsPE-FM to verity the effectiveness of prompt embedding, denoted as “w/o ${\bm{\hat{P}}}_{vi}$”. In the sixth setting, we omit the PEB from the Fusion Block, performing feature concatenation and convolution directly (denoted as “w/o FB-PEB”) to validate its role in fusion.
Table 1 and Figure 5 demonstrate the impact of each module on fusion performance. While removing any module leads to performance decrease, these changes may not be readily visible in the qualitative results but are clearly reflected in the quantitative data. As shown in Table 1, when infrared information is not used to supplement visible features, the metrics ${Q_{PI}}$ and ${Q_{NIQE}}$ increase significantly, and all fusion metrics decrease, confirming the effectiveness of incorporating infrared information to enhance visible features. If the HDE module is excluded, and infrared information is directly injected into the visible image, model performance does not improve. Instead, color distortion appears in the fusion results, and objective evaluation metrics decline to varying degrees. Similar issues are observed in Setting 3. Additionally, Setting 4 through 6 show slight degradation in detail retention, along with a decline in objective metrics, further validating the effectiveness of each module. Additional experiments, including hyperparameter analysis, model complexity comparisons, and discussions on method limitations and future work, are provided in the supplementary materials.
5 Conclusion

### v2

@@ IV-C Comparison with State-of-the-art Methods
In order to verify the effectiveness of our method, we compare it with two existing methods. The first methodology involves initially applying advanced image dehazing algorithms to remove haze from the visible images, followed by fusing the dehazed images with the infrared images. For this purpose, we select the latest and most effective dehazing methods, namely DIACMP [48] and Dehazeformer [49]. Next, we apply representative IR-VIS image fusion methods, such as MLFusion [33], U2Fusion [38], LRRNet [16], ALFusion [18], TIMFusion [50], and MRFS [51], to fuse the dehazed visible images with the infrared images. The second methodology employs the Text-IF method [7], which directly restores and fuses hazy images with the assistance of text information.
Experiments on MSRS dataset. To intuitively evaluate the fusion performance of different algorithms on the MSRS dataset, four pairs of IR-VIS images are selected, as shown in Fig. 3. As indicated from the red boxes in the first and second rows of Fig. 3, our method effectively preserves thermal radiation information, clearly highlighting the trousers of pedestrians, which most other methods fail to achieve. Although TIMFusion and Text-IF can also accomplish this to some extent, the results within the purple boxes reveal that they fail to accurately restore the texture details of trees, resulting in blurred outputs. In the red boxes of the third and fourth rows, our method preserves the details of windows while maintaining the scene brightness, producing a clear and well-restored window. In contrast, MLFusion, U2Fusion, and ALFusion suffer from brightness loss, leading to blurred scenes and poor visual effects. LRRNet, TIMFusion, MRFS, and Text-IF fail to deliver satisfactory contrast. Moreover, the other two sets of experimental results shown in Fig. 3 further highlight that the proposed method achieves superior visual performance compared to the competing methods.
Additionally, we conduct a quantitative comparison on 361 image pairs from the MSRS dataset to verify the effectiveness of our method. Tables I and II present the experimental results based on the DIACMP and Dehazeformer dehazing methods, respectively. As shown in Tables III, our method ranks first across all eight metrics, demonstrating its outstanding performance on the MSRS dataset. The higher $Q_{AB/F}$ and $Q_{CV}$ scores indicate that our fused images achieve superior detail clarity. Meanwhile, the best $Q_{SCD}$ and $Q_{VIF}$ scores suggest that our method ensures better visual consistency, closely matching the clear source images. The $Q_{MI}$ metric reflects the shared information between the fused and source images, confirming that our approach preserves more source image information. As perceptual quality metrics, $Q_{NIQE}$ and $Q_{PI}$ evaluate image naturalness and perceptual quality in a no-reference manner, where lower values indicate better alignment with human visual perception. Furthermore, $Q_{SF}$ measures the sharpness of image edges, demonstrating that our method produces haze-free fused images that are both natural and sharp.
Figure 3: Visual comparison of fusion results from different methods on the MSRS dataset. In the fusion results generated by the comparison methods, except for the last column, the first row of each image pair represents the result of first dehazing with DIACMP and then fusing. The second row shows the result of dehazing with Dehazeformer and then fusing. The last column represents the result of fusion using Text IF (first row) and our method (second row).
TABLE I: Quantitative analysis of our method compared with diacmp+fusion and text-if on the MSRS dataset. Red indicates the best, and blue indicates the second best.
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 1.706 0.274 608.391 1.121 0.199 5.022 4.285 9.953 U2Fusion 1.301 0.330 838.330 1.364 0.226 4.977 4.354 7.209 LRRNet 1.864 0.441 614.101 1.000 0.288 4.750 4.333 8.479 ALFusion 1.449 0.377 700.180 1.201 0.257 5.391 5.419 7.417 TIMFusion 1.800 0.382 1048.770 1.090 0.295 4.515 4.256 8.136 MRFS 1.580 0.457 332.068 1.348 0.325 5.068 4.800 9.042 Text-IF 1.428 0.558 458.252 1.351 0.369 4.260 4.312 8.393 Ours 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050
TABLE II: Quantitative analysis of our method compared with dehazeformer+fusion and text-if on the MSRS dataset. Red indicates the best, and blue indicates the second best.
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 1.720 0.270 613.291 1.119 0.199 5.052 4.337 9.826 U2Fusion 1.311 0.324 832.878 1.361 0.225 5.185 4.439 7.071 LRRNet 1.883 0.436 608.019 0.999 0.288 4.883 4.445 8.317 ALFusion 1.471 0.370 695.816 1.198 0.256 5.468 5.542 7.339 TIMFusion 1.785 0.382 1057.931 1.092 0.299 4.567 4.355 8.303 MRFS 1.603 0.448 328.901 1.340 0.323 5.150 4.905 8.843 Text-IF 1.428 0.558 458.252 1.351 0.369 4.260 4.312 8.393 Ours 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050
Figure 4: Visual comparison of fusion results from different methods on the M3FD dataset. In the fusion results generated by the comparison methods, except for the last column, the first row of each image pair represents the result of first dehazing with DIACMP and then fusing. The second row shows the result of dehazing with Dehazeformer and then fusing. The last column represents the result of fusion using Text IF (first row) and our method (second row).
TABLE III: Quantitative analysis of our method compared with diacmp+fusion and text-if on the M3FD dataset. Red indicates the best, and blue indicates the second best.
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 1.842 0.411 937.069 1.134 0.390 4.840 3.230 12.025 U2Fusion 1.607 0.511 805.163 1.266 0.358 4.908 3.377 12.009 LRRNet 1.579 0.473 724.325 1.195 0.358 4.355 3.245 11.529 ALFusion 1.510 0.419 890.521 1.205 0.322 4.878 3.960 9.291 TIMFusion 1.772 0.445 859.554 0.921 0.340 4.249 3.228 11.023 MRFS 1.756 0.514 666.466 1.208 0.415 4.304 3.277 13.071 Text-IF 1.935 0.542 634.406 1.189 0.427 5.277 3.902 11.940 Ours 2.070 0.571 640.661 1.357 0.440 4.144 3.158 14.140
TABLE IV: Quantitative analysis of our method compared with dehazeformer+fusion and text-if on the M3FD dataset. Red indicates the best, and blue indicates the second best.
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 1.879 0.400 1023.684 1.133 0.393 4.639 3.232 11.510 U2Fusion 1.647 0.497 824.810 1.281 0.355 5.251 3.635 11.295 LRRNet 1.592 0.463 737.595 1.206 0.360 4.528 3.455 11.116 ALFusion 1.503 0.403 922.165 1.227 0.324 5.095 4.196 8.974 TIMFusion 1.719 0.449 814.112 0.919 0.336 4.420 3.357 11.325 MRFS 1.788 0.509 693.528 1.200 0.417 4.467 3.460 12.551 Text-IF 1.935 0.542 634.406 1.189 0.427 5.277 3.902 11.940 Ours 2.070 0.571 640.661 1.357 0.440 4.144 3.158 14.140
Figure 5: Visual comparison of fusion results from different methods on the RoadScene dataset. In the fusion results generated by the comparison methods, except for the last column, the first row of each image pair represents the result of first dehazing with DIACMP and then fusing. The second row shows the result of dehazing with Dehazeformer and then fusing. The last column represents the result of fusion using Text IF (first row) and our method (second row).
TABLE V: Quantitative analysis of our method compared with diacmp+fusion and text-if on the RoadScene dataset. Red indicates the best, and blue indicates the second best.
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 2.350 0.457 509.388 1.248 0.411 3.703 3.196 12.083 U2Fusion 1.852 0.492 842.539 1.201 0.342 3.856 2.960 12.650 LRRNet 1.947 0.363 622.880 0.827 0.359 3.669 3.009 12.498 ALFusion 1.753 0.327 831.380 0.956 0.273 4.311 4.191 8.793 TIMFusion 2.349 0.356 699.810 0.881 0.425 4.484 4.241 10.307 MRFS 2.115 0.391 477.748 1.362 0.393 3.932 3.615 11.269 Text-IF 2.256 0.583 497.001 1.456 0.418 3.737 3.049 13.746 Ours 2.673 0.511 454.497 1.399 0.446 3.333 2.806 14.804
TABLE VI: Quantitative analysis of our method compared with dehazeformer+fusion and text-if on the RoadScene dataset. Red indicates the best, and blue indicates the second best.
Methods $Q_{MI}$$\uparrow$ $Q_{AB/F}$$\uparrow$ $Q_{CV}$$\downarrow$ $Q_{SCD}$$\uparrow$ $Q_{VIF}$$\uparrow$ $Q_{NIQE}$$\downarrow$ $Q_{PI}$$\downarrow$ $Q_{SF}$$\uparrow$ MLFusion 2.425 0.463 509.536 1.239 0.411 3.795 3.343 11.364 U2Fusion 1.864 0.495 862.820 1.175 0.340 3.873 3.029 12.117 LRRNet 1.948 0.360 617.051 0.780 0.357 3.534 2.967 12.326 ALFusion 1.726 0.315 844.762 0.934 0.257 4.321 4.258 8.545 TIMFusion 2.359 0.354 694.572 0.869 0.415 4.457 4.268 10.099 MRFS 2.119 0.396 472.542 1.332 0.389 3.847 3.619 10.922 Text-IF 2.256 0.583 497.001 1.456 0.418 3.737 3.049 13.746 Ours 2.673 0.511 454.497 1.399 0.446 3.333 2.806 14.804
Figure 6: Visualization of objective evaluation results. the group shows the results of using DIACMP for dehazing followed by fusion on MSRS, M3FD, and RoadScene.
Figure 7: Visualization of objective evaluation results. the group shows the results of using Dehazeformer for dehazing followed by fusion on MSRS, M3FD, and RoadScene.
Experiments on M3FD dataset. To evaluate the generalization ability of our method on the M3FD dataset, we select four pairs of IR-VIS images, with the visualization results shown in the Fig. 4. As illustrated in the first and second rows of Fig. 4, most methods exhibit blurring on the store signs, whereas our method preserves clear edges and texture details. Although MLFusion can also retain the texture information of the signs effectively, it performs poorly in restoring background details, resulting in background blurriness. This issue is also evident in the second and third image pairs, where MLFusion and other methods display color distortion in the sky regions. Thanks to the incorporation of infrared information during the dehazing stage in our method, the nextwork effectively restores these sky regions, achieving superior restoration and fusion results.
In terms of preserving thermal target information, our method also demonstrates a leading performance, which is particularly evident in the results shown in the fifth and sixth rows. Methods such as U2Fusion, LRRNet, ALFusion, and TIMFusion fail to retain thermal target information, leading to fusion results that do not effectively highlight thermal targets. In contrast, our method adopts a multi-stage prompt information injection strategy during the fusion phase, ensuring the infrared information is well-preserved. Although MLFusion, MRFS, and Text-IF can also emphasize targets to some extent, their performance in restoring background details remains suboptimal.
Quantitative comparison results on M3FD test set, based on the DIACMP and Dehazeformer dehazing methods, are presented in Tables III and IV, respectively. It can be observed that our proposed method ranks first in seven evaluation metrics and second in $Q_{CV}$, indicating its superior restoration and fusion capabilities.
Experiments on RoadScene dataset. We select four pairs of IR-VIS images from the RoadScene dataset to further evaluate the effectiveness and generalization capability of our method. The qualitative comparison results are presented in the Fig. 5. As shown in the first set of results in Fig. 5, our approach provides sharper object edges and more detailed textures. In the third and fourth rows, the content on the billboard within the red box is significantly clearer in our method compared to others, where varying degrees of blurriness are observed. Notably, the content displayed by Text-IF is completely unrecognizable.The results in the fifth and sixth rows indicate that our method can generate fused images with high contrast, preserving the original colors of the signboards while maintaining clear edges. From the results in the red box in the seventh and eighth rows, it can be seen that our method clearly highlights the information of distant vehicles, whereas the results generated by MLFusion, MRFS, and Text-IF are relatively blurry. Although other comparison methods can somewhat enhance vehicle information, as shown in the results in the purple box, their ability to restore the distant sky and roof areas is inferior to that of our method.
Quantitative comparison results on the RoadScene dataset are presented in Tables V and VI, which utilize the DIACMP and Dehazeformer dehazing methods respectively. Our method ranks first in six evaluation metrics, with $Q_{AB/F}$ and $Q_{SCD}$ ranking second. The best $Q_{NIQE}$ and $Q_{PI}$ scores indicate that our method generates fused images that are both natural and sharp.
Figure 8: Visualization of objective evaluation results. the group presents the objective evaluation results of fusion images obtained using DIACMP dehazing method, assessed with haze evaluation metrics.
Figure 9: Visualization of objective evaluation results. the group presents the objective evaluation results of fusion images obtained using Dehazeformer dehazing method, assessed with haze evaluation metrics.
To further validate the advantages of our method, we present violin plots for quantitative comparison in Figs. 69. These plots combine data density distribution (represented by the shape of the violin) with statistical features (embedded box plots). The width of each violin reflects data density, with wider sections indicating higher concentration. The embedded box plot shows key statistics: the red horizontal line indicates the mean, the box spans the 25% to 75% data range, the black line in the middle represents the median. The vertical axis represents the metric values, with the different-colored boxes corresponding to the results obtained by the various methods. These plots allow for a clear comparison of distribution, central tendencies, and variability across methods, highlighting performance differences effectively.
We categorize the metrics into two groups reflecting fusion quality and dehazing quality, organized based on the dehazing methods used. Fig. 6 presents violin plots of fusion metrics across three datasets, comparing our method with the combination of DIACMP for dehazing followed by fusion, and the Text-IF method. As shown in the $Q_{MI}$ metric’s violin plot, our method achieves the highest mean, with a more concentrated high-density distribution and relatively smaller variability, indicating that our data distribution is more centralized, demonstrating robust and stable performance in image fusion. Fig. 7 displays violin plots comparing our method with the combination of Dehazeformer for dehazing followed by fusion, and the Text-IF method. The results clearly show that our method exhibits superior fusion performance.
Fig. 8 illustrates violin plots of dehazing metrics across three datasets, comparing our method with the combination of DIACMP for dehazing followed by fusion, and the Text-IF method. As depicted in the $Q_{NIQE}$ metric’s violin plot, our method achieves the lowest score while showing a more concentrated distribution skewed towards lower values and exhibiting smaller variability compared to other methods. Fig. 9 shows violin plots comparing our method with the combination of Dehazeformer for dehazing followed by fusion, and the Text-IF method. The violin plots for all three metrics demonstrate that our method achieves stable and outstanding performance in image restoration.
Figure 10: Ablation study on the fusion network design. The first two columns are the input source images, and the third to ninth columns are different fusion network.
IV-D Ablation Study
We design six experimental settings to evaluate the effectiveness of each module. In the first setting, we remove the PGM-generated prompt ${\bm{\hat{P}}}_{ir}$, the haze density esitimation (HDE ) module, and the process of supplementing visible features with infrared features based on haze density, directly inputting ${\bm{F}}_{vi}$ into the Transformer Block for dehazing (denoted as “w/o ${\bm{F}}_{ir}$”) to assess the auxiliary role of infrared information. In the second setting, we replace the operation in Eq. (5) by directly adding ${\bm{\hat{F}}}_{ir}$ and ${\bm{F}}_{vi}$ (denoted as “w/o HDE”) to validate the impact of haze-density-based infrared integration. In the third setting, we remove the prompt embedding for ${\bm{\hat{P}}}_{ir}$ and the PEB module, injecting infrared features extracted by the encoder, guided by haze density ${\bm{H}}$, directly into ${\bm{F}}_{vi}$ (denoted as “w/o ${\bm{\hat{P}}}_{ir}$”) to evaluate the effect of prompt embedding. In the fourth setting, we test the effectiveness of the PEB module by removing it from IA-FRM, denoted as “w/o FR-PEB”. In the fifth setting, we remove ${\bm{\hat{P}}}_{vi}$ from MsPE-FM to verity the effectiveness of prompt embedding, denoted as “w/o ${\bm{\hat{P}}}_{vi}$”. In the sixth setting, we omit the PEB from the Fusion Block, performing feature concatenation and convolution directly (denoted as “w/o FB-PEB”) to validate its role in fusion.
Table VII and Fig. 10 demonstrate the impact of each module on fusion performance. While removing any module leads to performance decrease, these changes may not be readily visible in the qualitative results but are clearly reflected in the quantitative data. As shown in Table VII, when infrared information is not used to supplement visible features, the metrics ${Q_{PI}}$ and ${Q_{NIQE}}$ increase significantly, and all fusion metrics decrease, confirming the effectiveness of incorporating infrared information to enhance visible features. If the HDE module is excluded, and infrared information is directly injected into the visible image, model performance does not improve. Instead, color distortion appears in the fusion results, and objective evaluation metrics decline to varying degrees. Similar issues are observed in the setting “w/o ${\bm{\hat{P}}}_{ir}$”. Additionally, for the settings “w/o FR-PEB”, ‘w/o ${\bm{\hat{P}}}_{vi}$” and “w/o FB-PEB”, the experimental results show slight degradation in detail retention, along with a decline in objective metrics, further validating the effectiveness of each module.
@@ IV-D Ablation Study
TABLE VII: Quantitative results of six ablation experiments on the MSRS dataset, with the values in red indicating the best performance and those in blue indicating the second-best.
@@
IV-E Hyperparameters Analysis
A Our method involves two key hyperparameters: $\alpha$, used to balance the $L1$-loss, and $M$, representing the number of Fusion Blocks in MsPE-FM. During training, they are set to 2 and 5, respectively. To validate the rationale behind these choices, we conduct a detailed analysis of their impact on the model’s performance.
The impact of $\alpha$ on model performance. We fix $M$ at 5 and study the impact of different values of $\alpha$ within the range (0, 5] on the model performance. Table VIII presents the quantitative evaluation results of the fusion model on the MSRS dataset for various values of $\alpha$. The results indicate that when $\alpha$ is too small, the model performance does not reach its optimal level. Similarly, when $\alpha$ is too large, performance declines as well. The model achieves the best overall performance when $\alpha=2$, which validates the choice of $\alpha=2$ in this study.
The impact of $M$ on model performance. To analyze the impact of different values of $M$ on the model performance, we fix $\alpha$ at 2 and adjust $M$ within the range of [1, 7] to observe how the model performance changes. As shown in Table IX, the overall performance of the model improves with the number of Fusion Blocks. However, from the results for $M=5$ and $M=7$, it is evident that the performance gain has plateaued. Considering that increasing the number of Fusion Blocks significantly raises the model’s parameter count, we ultimately set $M$ to 5.
TABLE VIII: Quantitative analysis of six different $\alpha$ models on the MSRS dataset under the condition of $M=5$. Red indicates the best, and blue indicates the second best.
$\alpha$ $Q_{MI}$uparrow $Q_{AB/F}$uparrow $Q_{CV}$downarrow $Q_{SCD}$uparrow $Q_{VIF}$uparrow $Q_{NIQE}$downarrow $Q_{PI}$downarrow $Q_{SF}$uparrow 0.5 2.711 0.650 241.934 1.646 0.489 4.211 3.816 11.115 1 2.764 0.660 239.645 1.656 0.483 4.209 3.821 11.083 2 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050 3 2.743 0.658 243.385 1.650 0.488 4.105 3.819 11.049 4 2.787 0.656 240.587 1.657 0.489 4.149 3.838 11.054 5 2.030 0.372 387.993 1.572 0.345 5.321 5.401 7.059
TABLE IX: Quantitative analysis of four different $M$ models on the MSRS dataset under the condition of $\alpha=2$. Red indicates the best, and blue indicates the second best.
$M$ $Q_{MI}$uparrow $Q_{AB/F}$uparrow $Q_{CV}$downarrow $Q_{SCD}$uparrow $Q_{VIF}$uparrow $Q_{NIQE}$downarrow $Q_{PI}$downarrow $Q_{SF}$uparrow 1 2.567 0.651 251.174 1.669 0.484 4.113 3.816 10.967 3 2.915 0.662 244.317 1.643 0.488 4.146 3.839 10.982 5 2.720 0.652 238.420 1.662 0.490 4.110 3.811 11.050 7 2.832 0.659 238.883 1.656 0.491 4.108 3.840 11.027
IV-F Complexity Analysis
This paper employs a single-stage framework for hazy image fusion, which significantly reduces the model’s complexity and parameter count. To validate this advantage, we test the FLOPs and parameter count of each model and plot a bubble chart with the fusion metric $Q_{AB/F}$ to assess the model complexity. As shown in Fig. 11, our model achieves optimal performance while maintaining a low parameter count and the lowest computational complexity, fully demonstrating its efficiency and suitability for practical deployment.
Figure 11: Model Complexity Analysis. In the figure, the x-axis represents the FLOPs(G) for models with input images of size $256\times 256$, the y-axis denotes the average value of the fusion metric, and the radius of the bubbles reflects the model’s parameter count. In the method names, (1) refers to the result obtained by first applying DIACMP for dehazing followed by fusion, and (2) refers to the result obtained by first applying Dehazeformer for dehazing followed by fusion.
IV-G Limitations and Future Work
Although our method effectively handles hazy image fusion with high quality, real-world scenarios may still present challenging weather conditions such as low light, snow, and rain. Our approach does not currently account for the impact of these factors. While some existing methods have attempted to address this issue, they fail to effectively balance the impact of different degradations on the fusion results. Furthermore, these methods typically assume that the degradation types in the images have been encountered by the model, and the model performance tends to be suboptimal in scenarios with unseen degradations. Therefore, future work will focus on designing a multi-degradation joint processing framework to enable the model to effectively perform fusion and restoration even in scenarios with unseen degradations.
V Conclusion